Define covariant and contravariant tensors of rank two. What is the rank of a scalar
tensor? Show that the Kronecker delta function is a mixed tensor of rank 2
1
Expert's answer
2020-02-19T05:14:25-0500

Contravariant tensor of the second rank:-

If n 2 quantities A q s  in a coordinate system (x 1,x 2,........x n)are related to n 2 other quantities "\bar{A}"pr in another coordinate system("\bar{x^1}","\bar{x^2}" ,............"\bar{x^n}") by the transformation equations

"\bar{A}" pr= "\displaystyle\sum_{s=1}^n" "\displaystyle\sum_{q=1}^n" "{\partial\bar{x^p} \above{2pt}\partial x^q }" "{\partial\bar{x^r} \above{2pt}\partial x^s }" A q s p,r=1,2,..........,n

or, by our conventions,

"\bar{A}" pr="{\partial\bar{x^p} \above{2pt}\partial x^q }" "{\partial\bar{x^r} \above{2pt}\partial x^s }"

they are called components of a contravariant tensor of the second rank (or of rank two).

Covariant tensor of the second rank :-

If n 2 quantities A q s in a coordinate system (x 1,x 2,.......x n) are related to n 2 other quantities "\bar{A}" pr in another coordinate system ("\bar{x^1}",.........,"\bar{x^n}" ) by the transformation equations

"\bar{A}" pr= "\displaystyle\sum_{s=1}^n" "\displaystyle\sum_{q=1}^n" "{\partial x^q\above{2pt}\partial\bar{x^p} }" "{\partial x^s \above{2pt}\partial\bar{x^r} }" A q s p,r=1,2,........,n

or, by our conventions,

"\bar{A}" pr="{\partial x^q \above{2pt}\partial \bar{x^p} }" "{\partial x^s \above{2pt}\partial\bar{x^r} }"

they are called components of a covariant tensor of the second rank.

Show that the Kronecker delta function is a mixed tensor of rank 2

The Kronecker delta or Kronecker tensor, written, "\delta^p_q" and defined by

"\delta^p_q" ="\begin{cases}
 0 &\text{if } p{=}\mathllap{/\,}q \\
 1 &\text{if } p=q
\end{cases}"

is a mixed tensor of rank 2,justify the notation used.

Proof:-

we must prove that

"\bar{δ^j_k}" ="{\partial\bar{u^j} \above{2pt}\partial u^p }" "{\partial u^q \above{2pt} \partial\bar{u^k} }" "\delta^p_q"

By definition

"\bar{δ^p_q}=\delta^p_q=" "\begin{cases}
 0 &\text{if } p{=}\mathllap{/\,}q \\
 1 &\text{if } p=q
\end{cases}"

Coordinates "\bar{u ^j}" are functions of coordinates u p which are in turn functions of coordinates "\bar{u ^k}" .Then,by chain rule,

"{\partial \bar{u ^j} \above{2pt}\partial u^p }{\partial u^p \above{2pt}\partial \bar{u ^k} }={\partial \bar{u ^j} \above{2pt}\partial \bar{u ^k} }=\delta^j_k"

Now we must prove that the right side of the transformation equations of the mixed tensor above equals "\bar{\delta^j_k }" .It does:

"{\partial \bar{u ^j} \above{2pt}\partial u^p }{\partial u^q\above{2pt}\partial \bar{u^k} } \delta^p_q" ="{\partial \bar{u ^j} \above{2pt} \partial u^p}{\partial u^p \above{2pt}\partial \bar{u ^k} }" ="{\partial \bar{u^j} \above{2pt}\partial \bar{u^k} }=\delta^j_k=\bar{\delta^j_k}"

Rank of Scalar tensor:-

In fact tensors are merely a generalization of scalars and vectors; a scalar is a zero rank tensor, and a vector is a first rank tensor. The rank (or order) of a tensor is defined by the number of directions (and hence the dimensionality of the array) required to describe it.

A tensor is a mathematical representation of a scalar (tensor of rank zero), a vector(rank 1)

scalar=rank 0->magnitude ,no direction.

vector=rank 1->magnitude and direction.